If one of the solution of the equation x^3-2x^2+ax+10=0 is the additive inverse of another,then which inequality is true? a) a is between -40 and -30 b)a is between -30 and -20c)a is between -20 and -10d)a is between -10 and 0

Grade:11

1 Answers

sonika p
6 years ago
Answer is d, a is between -10 and 0
solution, let one root be alpha, then additive inverse is -alpha. let other root be beta. 
Now, alpha + (-alpha) + beta = -b/a = 2
thus beta =2
since beta is root and is equal to 2
thus 2^3-2*2^2+a*2+10=0 
8 – 8 +2a +10=0
2a = 10, a = -5
a is between -10 and 0